How do I solve the equation x/3-x-3/7=-1?: How do I solve the equation x/3-x-3/7=-1?
Answer by Alan3354(1431)

How do I solve the equation x/3-x-3/7=-1?
-------------
Multiply by 21 to get rid of fractions.
7x -21x-9 = -21
-14x = -12
x = 6/7
--------
If you meant x/3 - (x-3)/7 = -1
Mult by 21
7x - 3*(x-3) = -21
7x -3x +9 = -21
4x = -30
x = -7.5
Parentheses can eliminate ambiguities, and they're free.